In an alloy of 100 pounds of zinc and copper there are 75 pounds of copper. How much copper must be added so that the alloy may be 10% zinc?
Answer by stanbon(75887)

In an alloy of 100 pounds of zinc and copper there are 75 pounds of copper. How much copper must be added so that the alloy may be 10% zinc?
-------
copper + copper = copper
0.75*100 + 1*x = 0.90(100+x)
-----
Multiply thru by 100 to get:
75*100 + 100x = 90*100 + 90x
10x = 15*100
x = 150 lbs (amt of copper that must be added)
===================
Cheers,
Stan H.